OpenStudy (anonymous):

What is the exact value of cos(2 sin^-1 (1/3))

OpenStudy (loser66):

\(sin^{-1}(1/3)=?\)

OpenStudy (anonymous):

well don't you have to draw a triangle? 

OpenStudy (anonymous):

so 1^2 + b^2 = 3^2 so b=2sqrt2

OpenStudy (loser66):

then?? that is the side, not the angle.

OpenStudy (loser66):



OpenStudy (zarkon):

you don't need to find the angle

OpenStudy (anonymous):

ok thats what i thought

OpenStudy (loser66):

:)

OpenStudy (zarkon):

\[\cos(2\theta)=\cos^2(\theta)-\sin^2(\theta)\]

OpenStudy (zarkon):

then use your triangle

OpenStudy (anonymous):

and that is a trig identity?

OpenStudy (anonymous):

of a double angle

OpenStudy (zarkon):

yes it is a trig identity

OpenStudy (anonymous):

thanks

OpenStudy (zarkon):

can you get the answer now?

OpenStudy (anonymous):

yup it is 7/9

OpenStudy (anonymous):

thank you so much

OpenStudy (zarkon):

good